Plot a graph of cosx


1
Expert's answer
2021-11-02T19:04:20-0400
"f(x)=\cos x"

Domain: "(-\infin, \infin)"

Range: "[-1, 1]"

There are no asymptotes.


"f(-x)=\cos(-x)=\cos x=f(x), x\in \R"

The function is even. The graph is symmetric with respect to the y-axis.


"f(x+T)=f(x+2\pi)=f(x), x\in \R"

The function is periodic. The period is "T=2\pi."


"y" - intersection: "x=0, f(0)=\cos(0)=1."

Point "(0, 1)"


"x" - intersection(s): "y=0, 0=\cos x=>x=\dfrac{\pi}{2}+\pi n, n\in \Z."

Points "(\dfrac{\pi}{2}+\pi n, 0), n\in \Z"



"f'(x)=(\cos x)=-\sin x"

Find the critical number(s)


"f'(x)=0=>-\sin x=0=>x=\pi k, k\in \Z"

If "-\pi+2\pi k<x<2\pi k, k\in \Z," then "f'(x)>0, f(x)" increases.


If "2\pi k<x<\pi+2\pi k, k\in \Z," then "f'(x)<0, f(x)" decreases.


"f(2\pi k)=\cos(2\pi k)=1,k\in \Z"


"f(\pi+2\pi k)=\cos(\pi+2\pi k)=-1,k\in \Z"


The function "f(x)=\cos x" has the local maximum with value of "1" at "x=2\pi k, k\in \Z."


The function "f(x)=\cos x" has the local minimum with value of "-1" at "x=\pi+2\pi k, k\in \Z."



"f''(x)=(-\sin x)'=-\cos x"

Find the point(s) of inflection


"f''(x)=0=>-\cos x=0=>x=\dfrac{\pi}{2}+\pi m, m\in \Z""f(\dfrac{\pi}{2}+\pi m)=\cos(\dfrac{\pi}{2}+\pi m)=0,m\in \Z"

Points of inflection: "(\dfrac{\pi}{2}+\pi m, 0), m\in\Z"

If "-\dfrac{\pi}{2}+2\pi m<x<\dfrac{\pi}{2}+2\pi m, m\in \Z," then "f''(x)<0, f(x)" is concave down.


If "\dfrac{\pi}{2}+2\pi m<x<\dfrac{3\pi}{2}+2\pi m, m\in \Z," then "f''(x)>0, f(x)" is concave up.


Sketch the graph of "f(x)=\cos x"